If the base of a rectangle 2 1/3 cm and the perimeter is 11 7/15 cm, what is the height of the rectangle?
Answer by jorel1380(3719)

172/15-2(2 1/3)=172/15-14/3=
172/15-70/15=102/15 divided by 2=102/30=51/15
The height is 51/15 cm..